# Thermodynamic Limit in Learning Period Three

### v1

@@ Abstract
Figure 2 depicts how $f_{\infty}^{*}$ for $\phi=\text{erf}$ changes as $c$ varies. Note that the trained network always includes ${\cal D}$ since $f_{N}^{*}({\cal X})=f_{\infty}^{*}({\cal X})={\cal Y}$. In the case of ${\cal D}$ becoming the attractor ($-3\leq c<-2.2$ and $0<c\leq 0.22$), which corresponds to the successful attractor embedding by a learning machine, we observe that $f_{\infty}^{*}$ has no untrained attractors. However, varying $c$ changes the stability of all the periods—including ${\cal D}$—together, resulting in the emergence of untrained attractors: untrained period-3 ($-0.43\leq c<0$ and $2\leq c\leq 3$), chaotic ($c=-0.5,0.3$, etc.), and multiple attractors ($0.742\leq c\leq 0.764$, etc.). For $\phi=\sin,\cos,\text{ReLU}$, the bifurcation diagrams look very different from that in Fig. 2, but they have the same properties: only a handful of attractors appear at a time, which can be untrained ones (see Sec. LABEL:S:sec:dynamics in the Supplementary Material [23] for the bifurcation diagrams with the other network settings and the corresponding $\lambda_{T}$). We note that as the NTKs for $\phi=\text{erf},\sin,\cos,\text{ReLU}$ depend on the scaling and translation of inputs, varying $a$ or $b$, which we fixed, also yields a different bifurcation diagram. Nevertheless, the bifurcation structure has universal symmetry properties under the swapping of $a$ and $b$ (Sec. LABEL:S:sec:symmetry in the Supplementary Material [23]). In particular, there is a qualitative correspondence between the outside ($c<a,b<c$) and the inside ($a<c<b$) bifurcation structure, if $b\approx-a$ (Fig. 2(a)).
Figure 3: Two-dimensional slices of the characteristic bifurcation structures for $\phi=\text{erf}$ (top left), $\sin$ (top right), $\cos$ (bottom left), and ReLU (bottom right) with respect to $c$ and $\sigma$, with $a=-1$, $b=1$, $x_{0}=0$, and $T=10^{4}$. The black-hatched areas indicate the regions in which $\left|\frac{d}{dx}f_{\infty}^{*}(a)\frac{d}{dx}f_{\infty}^{*}(b)\frac{d}{dx}f_{\infty}^{*}(c)\right|<1$ holds. They correspond to the regions of ${\cal D}=\{a,b,c\}$ being locally stable, implying that for $\phi=\text{erf},\sin,\cos$, increasing $\sigma$ stabilizes ${\cal D}$. The blue-hatched area in the bottom-right panel indicates the region $(c,\sigma)$, where the trajectory heads toward infinity.
Figure 4: Two-dimensional slices of the characteristic bifurcation structure for $\phi=\text{erf}$ with respect to $c$ and $\sigma$, with $a=-1$, $b=1$, and $T=10^{4}$. (a) Number of kinds of period of attractors $N_{p}$ calculated with 100 equidistant initial states in $-10\leq x_{0}\leq 10$. $f_{\infty}^{*}$ has roughly $N_{p}$ attractors. (b) The zoom of (a). (c) The zoom of the top-left panel of Fig. 3. The black-hatched areas in (a),(b) indicate the regions of ${\cal D}=\{a,b,c\}$ being locally stable.
As we expect from the above discussion, we observe many types of characteristic attractors by varying $\sigma$ along with ${\cal D}$ (Figs. 3 and 4). Note that as $f_{\infty}^{*}$ may have multiple attractors (Fig. 4(a),(b)), certain trajectories reach attractors that are not depicted in Fig. 3. By utilizing these diagrams, we can control the untrained attractors of $f_{\infty}^{*}$ to obtain a large variety of attractors, which is extremely useful for generating intended dynamics. For example for $\phi=\text{erf}$, we obtain the period-13 attractor at $(c,\sigma)=(-0.665,1.32)$, period-19 attractor at $(c,\sigma)=(-0.655,1.172)$, and chaotic attractor at $(c,\sigma)=(-0.675,1.35)$ (“$+$” in Fig. 4(c)). For $\phi=\text{ReLU}$, the $\sigma$-dependence of NTK (Eq.(15)) is canceled out, resulting in the uniform bifurcation structure along the $\sigma$-direction (the bottom-right panel of Fig. 3). In contrast, for $\phi=\text{erf},\sin,\cos$, we can stabilize ${\cal D}$ by increasing $\sigma$ because the derivative of NTK at data point $y$ approaches zero for large $\sigma$ (the black-hatched areas in Fig. 3):
$\frac{\partial}{\partial x}\Theta^{\text{erf}}(y,y)=\frac{4\sigma^{2}y}{\pi\left[1+2\sigma^{2}(1+y^{2})\right]\sqrt{1+4\sigma^{2}(1+y^{2})}},$
$\frac{\partial}{\partial x}\Theta^{\sin}(y,y)=-\frac{\partial}{\partial x}\Theta^{\cos}(y,y)=\sigma^{2}ye^{-2\sigma^{2}(1+y^{2})}.$
However, $f_{\infty}^{*,\sin}$ and $f_{\infty}^{*,\cos}$ exponentially decrease to zero except at ${\cal D}$ as $\sigma$ grows; perturbations of $x_{0}$ from ${\cal D}$ lead to their different attractors for large $\sigma$; $\Theta^{\sin}(x,y)$ and $\Theta^{\cos}(x,y)$ become the Kronecker-delta-like functions in the infinite input-intensity limit $\sigma\rightarrow\infty$:
$\lim_{\sigma\rightarrow\infty}\Theta^{\sin}(x,y)=\lim_{\sigma\rightarrow\infty}\Theta^{\cos}(x,y)=\frac{1}{2}\bm{1}_{y}(x),$ (16)
where $\bm{1}_{y}(x)\equiv 1$ for $x=y$ and 00 otherwise. Interestingly, for $\phi=\text{erf}$, $\Theta^{\text{erf}}(x,y)$ and, therefore, $f_{\infty}^{*,\text{erf}}$ becomes piecewise-monotonic and piecewise-smooth, with ${\cal D}$ being its singular points in the limit $\sigma\rightarrow\infty$:
$\lim_{\sigma\rightarrow\infty}\Theta^{\text{erf}}(x,y)=\frac{2}{\pi}\arcsin\frac{1+xy}{\sqrt{(1+x^{2})(1+y^{2})}},$ (17)
$\frac{\partial}{\partial x}\lim_{\sigma\rightarrow\infty}\Theta^{\text{erf}}(x,y)=\frac{-2}{\pi(1+x^{2})}\cdot\frac{x-y}{|x-y|},$ (18)
enabling the candidates of robust chaos [41, 42] to appear. In addition, in this limit, NTK for sigmoidal activation (17) is equivalent to that for binary activation ($\phi=\text{sgn}$) [36]. Although $f_{N}^{*}$ for $\phi=\text{sgn}$ is beyond the scope of theorems 1 and 2, it asymptotically approaches a continous map $f_{\infty}^{*}$, exhibiting multiple stable periodic orbits for large $N$. Finite-size effects also emerge for $\phi=\sin,\cos$, which are caused by the wavy deviations of $f_{N}^{*}$ from $f_{\infty}^{*}$ (Sec. LABEL:S:sec:dynamics in the Supplementary Material [23]).
Finally, we discuss the peculiarity of learning period three with MNS and the scalability of our thermodynamic limit analysis in dynamical systems learning. Since $f_{\infty}^{*}(x)$ is described by the multiplication of $\Theta(x,{\cal X})$ and $\Theta^{-1}$, $|{\cal D}|$ crucially affects the complexity of the theoretical analysis of $f_{\infty}^{*}$. Accordingly, the longer the target orbit becomes—such as in the case of learning chaos and learning continuous-time dynamical systems from time series—the fewer the benefits of our proposed approach. However, similar to learning period three, if we adopt the strategy of exploiting the intrinsic mathematical structure of a learning machine with a small amount of training data, infinite-width networks provide a novel perspective to dynamical systems learning. We also investigated how the learned periods depend on the target period number $n$ and found that regardless of the choice of $\phi$, the number of unstable periods tends to increase dramatically after $n=3$ (Sec. LABEL:S:sec:LPn in the Supplementary Material [23]). This result suggests that a rich characteristic bifurcation structure is obtained by a combination of bulk unstable periods ensured by learning period three and a change in their stability through the system’s nonlinearity, which is our answer to the very first question.
@@
As an engineering application, our results suggest that various types of dynamics, including higher-order periodic orbits and chaos, can be obtained by teaching only three target data points to the network with appropriate settings. More importantly, the network may be a physical system or neuromorphic device, such as the physical reservoir computer [43]; embedding desirable dynamics into physical systems would be highly beneficial, for example, in a high-radiation environment in which a computer-simulated attractor may break down [44]. We leave the thermodynamic limit analysis of a physical neural network for future work.

### v2

@@ Abstract
Figure 2 depicts how $f_{\infty}^{*}$ for $\phi=\text{erf}$ changes as $c$ varies. Note that as long as $\hat{\Theta}$ and $\Theta$ have full rank, trained networks completely learn the target orbit ${\cal D}$ since $f_{N}^{*}({\cal X})=f_{\infty}^{*}({\cal X})={\cal Y}$. In the case of ${\cal D}$ becoming the attractor ($-3\leq c<-2.2$ and $0<c\leq 0.22$), which corresponds to the successful attractor embedding by a learning machine, we observe that $f_{\infty}^{*}$ has no untrained attractors. However, varying $c$ causes the bifurcation of the embeddable attractors of $f_{\infty}^{*}$, resulting in the emergence of untrained attractors: untrained period-3 ($-0.43\leq c<0$ and $2\leq c\leq 3$), chaotic ($c=-0.5,0.3$, etc.), and multiple attractors ($0.742\leq c\leq 0.764$, etc.). Consequently, only a handful of attractors appear at a time, and almost all periods latently exist as unstable periods (Fig. 2(a),(b)). Such bifurcation of the characteristic attractors strongly depends on the network settings (see Sec. LABEL:S:sec:dynamics in the SM [26] for the bifurcation diagrams with respect to $c$ and $\sigma$, along with the corresponding $\lambda_{T}$, which includes Refs. [43, 44]). We note that as the NTKs for $\phi=\text{erf},\sin,\cos,\text{ReLU}$ depend on the scaling and translation of inputs, varying $a$ or $b$, which we fixed, also yields a different bifurcation. Nevertheless, the bifurcation structure has universal symmetry properties under the swapping of $a$ and $b$ (Sec. LABEL:S:sec:symmetry in the SM [26]). In particular, if $\Theta(x,y)=\Theta(-x,-y)$, $\left.f_{\infty}^{*}\right|_{{\cal D}=\{a,-a,c\}}$ is topologically conjugate to $\left.f_{\infty}^{*}\right|_{{\cal D}=\{-a,a,-c\}}$, and there is a qualitative correspondence between the outside ($c<a,b<c$) and the inside ($a<c<b$) bifurcation structure, if $b\approx-a$ (Fig. 2(a),(b)).
Figure 3: Change in dynamical system $\sigma_{\text{fb}}f_{\infty}^{*}$ for $\phi=\text{erf}$ with respect to feedback strength $\sigma_{\text{fb}}$, with ${\cal D}=\{-1,1,0\}$ and $\sigma=1$. (a) Bifurcation diagram calculated with $-10\leq x_{0}\leq 10$ and $T=10^{5}$. (b) Stability changes in the acquired periods $\{x_{p}\}$ of period $p$, calculated by $\lambda_{p}\equiv\ln\left|\frac{d}{dx}(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})\right|$ using the solutions of $(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})=x_{p}$. The vertical dotted lines correspond to the network state in LP3 ($\sigma_{\text{fb}}=1$). The horizontal dotted line indicates the boundary of the stability $\lambda_{p}=0$; $\lambda_{p}<0$ and $\lambda_{p}>0$ mean (locally) stable and unstable, respectively. In this setting, decreasing $\sigma_{\text{fb}}$ toward 0 externalizes the latently existing unstable periods as period-doubling cascades, which is illustrated as the continuous lines of $\{\lambda_{p}\}$ from $\lambda_{p}>0$ (at $\sigma_{\text{fb}}=1$) to $\lambda_{p}<0$ (at some $\sigma_{\text{fb}}\in(0,1)$).
So far, we have only focused on the characteristic attractors of the neural networks in LP3, which are just small parts of latently existing infinitly many periodic orbits. To illuminate the meaning of the latently existing unstable periods, we further extend our learning procedure to varying the readout weights of the trained neural networks [25]. Figure 3 depicts how its attractor changes as the scale of the readout weights—the feedback strength $\sigma_{\text{fb}}$—varies. We set the network state of LP3 to $\sigma_{\text{fb}}=1$; this dynamical system is expressed as $\sigma_{\text{fb}}f_{\infty}^{*}$. Since the network is prohibited from having any periodic orbits except a fixed point $x=0$ at $\sigma_{\text{fb}}=0$, all the acquired periods $\{x_{p}\}$ must disappear at some $\sigma_{\text{fb}}\in(0,1)$ with $\frac{d}{dx}(\sigma_{\text{fb}}f_{\infty}^{*})^{p}(x_{p})=1$ (tangent bifurcation) due to the differentiability of $f_{\infty}^{*}$; some of them may become attractors before dying out. The way of the externalization of latently unstable periods depends on the choice of ${\cal D}$ and the network structure (see Fig. LABEL:S:fig:externalize_inf in the SM [26]).
Now, let us assume that ${\cal D}=\{a,-a,0\}$ (symmetric LP3) and $\phi=\text{erf},\sin,\cos$. Surprisingly, this symmetry of ${\cal D}$ and differentiability of $\phi$ will induce further universal bifurcation that externalizes all the latently unstable periods into stable ones after learning, corresponding to the Sharkovsky ordering (Eq. (1)) (see Figs. LABEL:S:fig:externalize_universal_infLABEL:S:fig:externalize_universal in the SM [26] for $\phi=\text{ReLU}$, which induces slightly different externalization). Thanks to the symmetry in the distribution of $W^{\text{in}}$ at $W^{\text{in}}=0$, $\Theta(x,x),\Theta(x,-x)$, and $\Theta(x,0)$ are even functions of $x$ since $\Theta(x,y)=\Theta(-x,-y)$. Applying $b=-a$ and $c=0$ to Eq. (9) and exploiting these symmetries, the derivative of $f_{\infty}^{*}$ at $x=0$ is given as follows:
$\left.\frac{df_{\infty}^{*}}{dx}(0)\right|_{{\cal D}=\{a,-a,0\}}=-\frac{a\frac{\partial\Theta}{\partial x}(0,a)}{\Theta(a,a)-\Theta(a,-a)}$ (14)
$=-\frac{\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)+{\cal O}(a)}{2\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)+{\cal O}(a)}\rightarrow-\frac{1}{2}\quad(a\rightarrow 0).$
Note that $\frac{\partial^{2}\Theta}{\partial x\partial y}(0,0)=\mathbb{E}_{\omega,\beta}\left[\omega^{2}\phi^{\prime}(\beta)^{2}\right]>0$. Next, for $a\ll 1$, the local shape of $f_{\infty}^{*}(x)$ around ${\cal D}=\{a,-a,0\}$ is well described as a polynomial function $g(x)$. Imposing $g({\cal X})={\cal Y}$ and $\frac{dg}{dx}(0)=-\frac{1}{2}$, its third-order approximation is uniquely determined as
$g(x)=a\hat{g}\left(\frac{x}{a}\right),\quad\hat{g}(x)\equiv 1-\frac{x}{2}-\frac{3x^{2}}{2}+0\cdot x^{3},$ (15)
which is topologically conjugate to $\hat{g}$. Furthermore, under $\mu>2$ and $\sigma_{\text{fb}}>0$, $x_{n+1}=\sigma_{\text{fb}}\hat{g}(x_{n})$ is topologically conjugate to the logistic map $y_{n+1}=\mu y_{n}(1-y_{n})$ in the following manner [45]:
$x_{n}=-\frac{2}{3}\left[\frac{\mu}{\sigma_{\text{fb}}(\mu)}\left(-y_{n}+\frac{1}{2}\right)+\frac{1}{4}\right],$ (16)
$\sigma_{\text{fb}}(\mu)=\frac{4}{5}\sqrt{\frac{\mu}{2}\left(\frac{\mu}{2}-1\right)+\frac{1}{100}}-\frac{2}{25}.$ (17)
With $0<\sigma_{\text{fb}}(\mu_{1})\leq\sigma_{\text{fb}}(\mu_{2})$ for $2<\mu_{1}\leq\mu_{2}$ and the Sharkovsky ordering in $\mu$ [46, 47], we obtain the Sharkovsky ordering in externalization:
$\sigma_{\text{fb}}[3]$ $\geq\sigma_{\text{fb}}[5]\geq\cdots\geq\sigma_{\text{fb}}[2\cdot 3]\geq\sigma_{\text{fb}}[2\cdot 5]\geq\cdots$ (18)
$\geq\sigma_{\text{fb}}[2^{2}\cdot 3]\geq\cdots\geq\sigma_{\text{fb}}[2^{2}]\geq\sigma_{\text{fb}}[2]\geq\sigma_{\text{fb}}[1],$
where we denote by $\sigma_{\text{fb}}[n]$ the least value of $\sigma_{\text{fb}}$ for which $\sigma_{\text{fb}}\hat{g}$ possesses period $n$ as its attractor and $\sigma_{\text{fb}}[1]\equiv 0$. Here, $10=\sigma_{\text{fb}}[1]\leq\cdots\leq\sigma_{\text{fb}}[3]=\sigma_{\text{fb}}(1+\sqrt{8})<$ 1. Consequently, the universality (Eqs. (15) and (18)) retains no information about network settings $N,\phi,\sigma$ or the dataset $a$, implying the strong convergence of the externalization in symmetric LP3 (see Sec. LABEL:S:sec:externalize in the SM [26] for a detailed discussion, which includes Ref. [48]). Interestingly, with a small $a$, we observed the coexistence of the controlled attractors—which come from the universality $\hat{g}$—and the characteristic attractors—which come from the details of the network structure—in the externalization $\sigma_{\text{fb}}f_{\infty}^{*}$ (see Fig. LABEL:S:fig:externalize_universal_inf in the SM [26]).
Although LP3 is not a necessary condition for achieving period three (learning period two or even random neural networks [49] may have period three, see Fig. LABEL:S:fig:period_n_distribution and Fig. LABEL:S:fig:randomNN in the SM [26]), it provides a sufficient condition for embedding attractors with all periods as a universal bifurcation along with externally controllable parameters ${\cal D}$ and $\sigma_{\text{fb}}$—which is our answer to the very first question. LP3 also provides new yet important perspectives in the learning of dynamics. First, even if neural networks completely learn a target orbit ${\cal D}$, they may fail to replicate ${\cal D}$, since its stability depends on the local structure of the trained map $f_{\infty}^{*}$ around ${\cal D}$. Second, LP3 is not a goal, but a groundwork or a primer for updating the connectivity of random networks to generate all types of periodic orbits, including chaos, after learning. Generating network dynamics with a minimal dataset is not just efficient but also compatible with the theoretical analysis, since $f_{\infty}^{*}(x)$ is described by the multiplication of $\Theta(x,{\cal X})$ and $\Theta^{-1}$; learning dynamics from time series lead to a large $|{\cal D}|$ and a nearly singular $\Theta$, making the analysis of $f_{\infty}^{*}$ ineffective.
@@
As an engineering application, we may replace the feedforward network part with a physical system or neuromorphic device [50, 51, 52]; embedding desirable dynamics into physical systems would be highly beneficial, for example, in a high-radiation environment in which a computer-simulated attractor may break down [53]. We leave the further analysis on the control of embeddable attractors in physical systems for future work.